OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x – y = –94x – 3y = –22 a) 4(–8x – 9) – 3y = –22 B) 4x – 3(–8x – 9) = –22 c) 4(8x + 9) – 3y = –22 d) 4x – 3(8x + 9) = –22

OpenStudy (turingtest):

are the two equations 8x – y = –94 and x – 3y = –22 ? because it's not so clear when you copy-paste like that ;)

OpenStudy (anonymous):

Sorry lol but they are two seperate equations like you typed.

OpenStudy (turingtest):

ok, first solve the first equation for y what do you get?

OpenStudy (anonymous):

y=8x+94

OpenStudy (turingtest):

typo put that expression for y where the y is in the next expression

OpenStudy (anonymous):



OpenStudy (anonymous):

and when i foil do i multiply by x or 3

OpenStudy (turingtest):

yes that is the right expression and no you don't FOIL, you distribute the 3

OpenStudy (anonymous):

I don't understand what you mean to distribute?

OpenStudy (turingtest):

foil is used when multiplying two binomials\[(a+b)(c+d)=ac+ad+bc+bd\]distribution is just\[a(b+c)=ab+ac\]so "distribute" the 3 to the parentheses

OpenStudy (anonymous):

so i would go 8 (3) and 94(3)

OpenStudy (turingtest):

first make sure you are careful with the signs distribute -3, not 3 and just don't just forget to leave the x next to -3(8)x other than that though, yeah, right idea

OpenStudy (anonymous):

oh alright thank you for your help i was able to solve some other problems too!!!!